# A single-loop SPIDER-type stochastic subgradient method for expectation-constrained nonconvex nonsmooth optimization

### v1

@@ Abstract
Many real-world problems, such as those with fairness constraints, involve complex expectation constraints and large datasets, necessitating the design of efficient stochastic methods to solve them. Most existing research focuses on cases with no constraint or easy-to-project constraints or deterministic constraints. In this paper, we consider nonconvex nonsmooth stochastic optimization problems with expectation constraints, for which we build a novel exact penalty model. We first show the relationship between the penalty model and the original problem. Then on solving the penalty problem, we present a single-loop SPIDER-type stochastic subgradient method, which utilizes the subgradients of both the objective and constraint functions, as well as the constraint function value at each iteration. Under certain regularity conditions (weaker than Slater-type constraint qualification or strong feasibility assumed in existing works), we establish an iteration complexity result of $O(\epsilon^{-4})$ to reach a near-$\epsilon$ stationary point of the penalized problem in expectation, matching the lower bound for such tasks. Building on the exact penalization, an $(\epsilon,\epsilon)$-KKT point of the original problem is obtained. For a few scenarios, our complexity of either the objective sample subgradient or the constraint sample function values can be lower than the state-of-the-art results in [9, 18, 30] by a factor of $\epsilon^{-2}$. Moreover, on solving two fairness-constrained problems, our method is significantly (up to 466 times) faster than the state-of-the-art algorithms, including switching subgradient method in [18] and inexact proximal point methods in [9, 30].
@@
stochastic, subgradient, expectation constraints, weakly convex, fairness constrained classification.
@@ 1 Introduction
Stochastic optimization is fundamental to the statistical sciences and serves as the backbone of data-driven learning processes. In this paper, we study this topic by designing first-order methods (FOMs) for solving the expectation-constrained nonconvex programming:
(P) $\min_{{\mathbf{x}}}\,\,{f({\mathbf{x}})},\text{ s.t. }\underbrace{\mathbb{E}_{\xi_{g}\sim\mathbb{P}_{g}}[g({\mathbf{x}},\xi_{g})]}_{=:g({\mathbf{x}})}\leq 0.$
Here, $\mathbb{P}_{g}$ is an unknown probability distribution, and we will consider a general distribution and a special uniform distribution on a finite set, while $f$ is always assumed to be stochastic. Notice that a problem with multiple constraints $g_{i}({\mathbf{x}})\leq 0,i=1,\ldots,m$ can be formulated into Eq. P by setting $g({\mathbf{x}})\equiv\max_{i=1,\ldots,m}g_{i}({\mathbf{x}})$.
Substantial research has been devoted to FOMs for solving function-constrained optimization problems. Among these FOMs, most are developed for deterministic function-constrained problems or for expectation-constrained convex optimization. This is well-documented in the literature for convex scenarios [6, 7, 21, 22, 23, 26, 29, 40, 41, 44] and nonconvex scenarios [3, 4, 10, 14, 28, 27, 31]. Only a few works (e.g., [2, 9, 18, 24]) have studied FOMs for expectation-constrained nonconvex optimization. We focus on nonconvex nonsmooth expectation-constrained problems in the form of Eq. P. Throughout this paper, we assume the follows.
@@
$\mathrm{(a)}$ The set ${\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)\cap\Omega$ is nonempty, where $\Omega:=\{{\mathbf{x}}:g({\mathbf{x}})\leq 0\}$, and ${\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$ is bounded. Specifically, there exists $D>0$ such that $\|{\mathbf{x}}_{1}-{\mathbf{x}}_{2}\|\leq D$ for any ${\mathbf{x}}_{1},{\mathbf{x}}_{2}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$.
$\mathrm{(b)}$ The function $f$ is $\rho_{f}$-weakly convex, and $g$ is $\rho_{g}$-weakly convex for some $\rho_{f}>0$ and $\rho_{g}\geq 0$. That is, $f+\frac{\rho_{f}}{2}\|\cdot\|^{2}$ and $g+\frac{\rho_{g}}{2}\|\cdot\|^{2}$ are convex.
$\mathrm{(c)}$ The subdifferential sets of $f$ and $g$ are nonempty in their respective domains. In addition, it holds that $\|{\boldsymbol{\zeta}}_{f}\|\leq l_{f}$ and $\|{\boldsymbol{\zeta}}_{g}\|\leq l_{g}$ for all ${\boldsymbol{\zeta}}_{f}\in\partial f({\mathbf{x}})$ and ${\boldsymbol{\zeta}}_{g}\in\partial g({\mathbf{x}})$ for any ${\mathbf{x}}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$.
From items (a) and (c) in the above assumption, we have $|f({\mathbf{x}}_{1})-f({\mathbf{x}}_{2})|\leq l_{f}D=:B_{f}$ for any ${\mathbf{x}}_{1},{\mathbf{x}}_{2}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$. In item (b), we allow $\rho_{g}=0$, which indicates the convexity of $g$. However, we always assume $\rho_{f}>0$ and thus consider a nonconvex objective function $f$.
@@
OGC and CGC denote the number of calls to the sample subgradients of $f$ and $g$, respectively, while CFC represents the number of evaluations of the sample constraint function $g(\cdot,\xi_{g})$.
@@
The nonsmooth expectation constraints in Eq. P pose significant challenges for designing efficient and reliable stochastic methods, as the projection onto an expectation-constrained set is prohibitively expensive. To circumvent the costly projection step, several methods such as online convex method [43], augmented Lagrangian method (ALM) [24], inexact proximal point method (IPP) [9, 30] and switching subgradient method (SSG) [18, 23] have been explored. Among these, the online convex method, ALM, and IPP require at least two loops. The online convex method [43], developed under a convex setting, relies on a restrictive assumption that the noise associated with stochastic constraints is bounded. Under a regularity condition, [24] and [39] demonstrate the effectiveness of the ALM applied to problem Eq. P, yielding competitive numerical results. [39] establishes convergence to a KKT point. To derive complexity results to an $(\epsilon,\epsilon)$-KKT point (see Definition 1.2), [24] assumes additional smoothness on $g$ and exploits this structure to efficiently solve each nonconvex ALM subproblem by an accelerated method. Under a certain constraint qualification (CQ), [9] and [30] respectively provide convergence results and complexity analysis for stochastic IPP methods to find an $(\epsilon,\epsilon)$-KKT point. These methods iteratively solve strongly convex subproblems by incorporating a proximal term into the objective and constraint functions. Notably, the complexity results in [9] represent the state-of-the-art results for problem Eq. P under nonsmooth weakly convex settings.
@@
Double (or more)-loop algorithms can utilize optimal subroutines for solving inner subproblems to achieve low-order complexity results. However, implementing them needs extra efforts to design appropriate checkable stopping criteria for inner subsolvers or requires careful tuning of the inner iteration counts to reach a satisfactory empirical performance. This disadvantage is especially pronounced in stochastic settings, where the randomness complicates the verification of a stopping condition and efficient tuning. Hence, single-loop methods for solving expectation-constrained problems are preferred in practice. However, such methods are still limited. Both of [2, 36] design FOMs based on penalty or ALM for constrained stochastic programming but they require smoothness of the objective (possibly with a proximable nonsmooth convex term) and constraint functions. The authors in [18] apply a single-loop SSG [34] to problem Eq. P. The SSG is easier to implement while delivering competitive numerical performance. Under a (uniform) Slater-type CQ (see 5), the SSG adapts its stepsize depending on whether the current iterate is near feasible or not, switching between fixed and adaptive steps based on Polyak’s rules [34]. This adaptation is crucial in maintaining near feasibility to achieve a near-KKT point. One critical drawback of the SSG is the requirement of frequent checks for near feasibility at generated iterates, leading to a higher CFC and overall complexity compared to IPP in [9] for solving the nonconvex problem Eq. P. Moreover, in the stochastic setting, [18] requires a light-tail assumption.
@@
Without the need for feasibility checking, we propose a novel single-loop stochastic FOM for solving Eq. P, named 3S-Econ. At the $k$-th iteration, 3S-Econ requires unbiased stochastic subgradient $\boldsymbol{\zeta}_{f}^{(k)}$ of $f$ and $\boldsymbol{\zeta}_{g}^{(k)}$ of $g$, evaluated at the iterate ${\mathbf{x}}^{(k)}$, along with an approximation $u^{(k)}$ for $g({\mathbf{x}}^{(k)})$. Using this information, it performs the following update step:
@@
(1) ${\mathbf{x}}^{(k+1)}={\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{f}^{(k)}+\beta\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)\boldsymbol{\zeta}_{g}^{(k)}\right),$
@@
where $\beta>0$, $\nu>0$, and $\alpha_{k}>0$ denotes the step size at iteration $k$. We will show that 3S-Econ not only is easy to implement but also enjoys lower-order complexity than the best-known results to produce an $(\epsilon,\epsilon)$-KKT point of problem Eq. P under milder conditions; see Table 1.
@@ 1.1 Applications
Neyman-Pearson classification. Let $\phi(\mathbf{x};\mathbf{a},b)$ be a loss function for a binary classification problem. Here, $\mathbf{x}$ is the variable, and ${\mathbf{a}}$ is a data point with a label $b\in\{+1,-1\}$. Denote $c>0$ as the threshold of the false-positive error, $\left\{\mathbf{a}_{i}^{+}\right\}_{i=1}^{n^{+}}$ and $\left\{\mathbf{a}_{i}^{-}\right\}_{i=1}^{n^{-}}$ as the positive-class samples (i.e., label $=1$) and negative-class samples (i.e., label $=-1$) in the training dataset, respectively. The Neyman-Pearson classification aims to minimize the more concerned false-negative error while controlling false-positive error at an acceptable level by solving
(2) $\min_{\mathbf{x}\in\mathbb{R}^{d}}\frac{1}{n^{+}}\sum_{i=1}^{n^{+}}\phi\left(\mathbf{x};\mathbf{a}_{i}^{+},1\right),\text{s.t. }\frac{1}{n^{-}}\sum_{i=1}^{n^{-}}\phi\left(\mathbf{x};\mathbf{a}_{i}^{-},-1\right)-{c}\leq 0.$
Clearly, Eq. 2 is a special instance of Eq. P with $f$ having a finite-sum structure, and $\mathbb{P}_{g}$ being a uniform distribution on the negative-class samples.
@@
Fairness-constrained classification. This application aims to ensure that classification models treat different groups equitably, avoiding biased outcomes. Traditional models focus on maximizing accuracy, which can lead to unfair performance across different groups of data, such as males and females. Fairness constraints are applied to balance outcomes for all groups, using principles like demographic parity [1] or equalized odds [17]. Achieving fairness requires trade-offs, as improving fairness can reduce overall accuracy. This research area is particularly important in fields like healthcare and criminal justice [11]. Let us consider a classification problem with receiver operating characteristic (ROC)-based fairness [37]. Suppose there are three groups of data: an unlabeled set $\mathcal{D}=\left\{\left(\mathbf{a}_{i},b_{i}\right)\right\}_{i=1}^{n}$, a protected set $\mathcal{D}_{\mathrm{p}}=\left\{\left(\mathbf{a}^{\mathrm{p}}_{i},b^{\mathrm{p}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{p}}}$, and an unprotected set $\mathcal{D}_{\mathrm{u}}=\left\{\left(\mathbf{a}^{\mathrm{u}}_{i},b^{\mathrm{u}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{u}}}$. Here, the protected set contains the features associated with the minority group, while the unprotected set includes the features from the majority group. The fairness-constrained problem can be formulated as
(3) $\vspace{-1mm}\min_{{\mathbf{x}}:\|{\mathbf{x}}\|\leq D}\Phi_{\Theta}({\mathbf{x}}),\,\,\text{{s.t.}}\,\,\Phi({\mathbf{x}})\leq\Phi^{*}+\kappa_{1},$

### v2

# A SPIDER-type stochastic subgradient method for expectation-constrained nonconvex nonsmooth optimization

@@ Abstract
Many real-world problems, such as those with fairness constraints, involve complex expectation constraints and large datasets, necessitating the design of efficient stochastic methods to solve them. Most existing research focuses on cases, which have no constraint, or have constraints that admit an easy projection, or only have deterministic constraints. In this paper, we consider weakly convex stochastic optimization problems with expectation constraints, for which we build an exact penalty model. We first show the relationship between the penalty model and the original problem. Then on solving the penalty problem, we present a SPIDER-type stochastic subgradient method, which utilizes the subgradients of both the objective and constraint functions, as well as the constraint function value at each iteration. Under the Slater-type constraint qualification (CQ), we establish an iteration complexity result of $O(\epsilon^{-4})$ to reach a near-$\epsilon$ stationary point of the penalized problem in expectation, matching the lower bound for such tasks. Building on the exact penalization, an $(\epsilon,\epsilon)$-KKT point of the original problem is obtained. For a few scenarios, our complexity of either the objective sample subgradient or the constraint sample function values can be lower than the state-of-the-art results in [10, 27, 46] by a factor of $\epsilon^{-2}$. Moreover, on solving two fairness-constrained problems and a multi-class Neyman-Pearson classification problem, our method is significantly (up to 466 times in terms of data pass) faster than the state-of-the-art algorithms, including the switching subgradient method in [27] and the inexact proximal point methods in [10, 46].
@@
stochastic, subgradient, expectation constraints, weakly convex, fairness-constrained classification.
@@ 1 Introduction
Stochastic optimization is fundamental to the statistical learning and serves as the backbone of data-driven learning processes. In this paper, we study this topic by designing first-order methods (FOMs) for solving the expectation-constrained nonconvex programming:
(P) $\min_{{\mathbf{x}}\in\mathcal{X}}\,\,{f({\mathbf{x}})},\text{ s.t. }\underbrace{\mathbb{E}_{\xi_{{\mathbf{g}}}\sim\mathbb{P}_{{\mathbf{g}}}}[{\mathbf{g}}({\mathbf{x}},\xi_{{\mathbf{g}}})]}_{=:{\mathbf{g}}({\mathbf{x}})}\leq 0.$
Here, $\mathcal{X}\subset\mathbb{R}^{d}$ is a closed convex set that allows a computationally easy projection operator, ${\mathbf{g}}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ is a vector-valued function consisting of $m$ components $g_{1},\ldots,g_{m}$, and $\mathbb{P}_{{\mathbf{g}}}$ is an unknown distribution from which we can draw independent samples. We will consider both general distributions and the special case where $\mathbb{P}_{{\mathbf{g}}}$ is uniform over a finite sample set; for the latter we can obtain deterministic subgradient of each $g_{i}$ if needed. For $f$, we assume only its stochastic subgradient is available.
Substantial research has been devoted to FOMs for solving function-constrained optimization problems. Among these FOMs, most are developed for deterministic function-constrained problems or for expectation-constrained convex optimization. This is well-documented in the literature for convex scenarios [7, 8, 31, 32, 33, 40, 45, 59, 60, 64] and nonconvex scenarios [3, 4, 13, 19, 22, 25, 28, 42, 43, 47, 54]. Only a few works (e.g., [2, 10, 27, 38]) have studied FOMs for expectation-constrained nonconvex optimization. We focus on nonconvex nonsmooth expectation-constrained problems in the form of Eq. P. Throughout this paper, we make the following assumption.
@@
$\mathrm{(a)}$ The set $\mathcal{X}\cap\{{\mathbf{x}}:{\mathbf{g}}({\mathbf{x}})\leq 0\}$ is nonempty, and $\mathcal{X}$ is bounded. Specifically, there exists $D>0$ such that $\|{\mathbf{x}}_{1}-{\mathbf{x}}_{2}\|\leq D$ for any ${\mathbf{x}}_{1},{\mathbf{x}}_{2}\in\mathcal{X}$.
$\mathrm{(b)}$ The function $f$ is $\rho_{f}$-weakly convex on $\mathcal{X}$}, and $g_{i}$ is $\rho_{g}$-weakly convex on $\mathcal{X}$} for some $\rho_{f}\geq 0$, $\rho_{g}\geq 0$ and each $i\in[m]:=\{1,\ldots,m\}$. That is, $f+\frac{\rho_{f}}{2}\|\cdot\|^{2}$ and $g_{i}+\frac{\rho_{g}}{2}\|\cdot\|^{2}$ are convex on $\mathcal{X}$} for each $i\in[m]$.
$\mathrm{(c)}$ The subdifferential sets of $f$ and $g_{i}$ are nonempty in their respective domains for each $i\in[m]$. In addition, it holds that $\|{\boldsymbol{\zeta}}_{f}\|\leq l_{f}$ and $\|\boldsymbol{{\boldsymbol{\zeta}}}_{g_{i}}\|\leq l_{g}$ for all ${\boldsymbol{\zeta}}_{f}\in\partial f({\mathbf{x}})$ and $\boldsymbol{{\boldsymbol{\zeta}}}_{g_{i}}\in\partial g_{i}({\mathbf{x}})$ for any ${\mathbf{x}}\in\mathcal{X}$ and any $i\in[m]$.
From items (a) and (c) in the above assumption, we have
[…]
In item (c), $\partial$ refers to the subdifferential, see Definition 1.1. Item (c) holds naturally when $f$ and $\{g_{i}\}$ are all Lipschitz continuous.
@@
OGC and CGC denote the number of calls to the sample subgradients of $f$ and $\mathbf{g}$}, respectively, while CFC represents the number of evaluations of the sample constraint function ${\mathbf{g}}(\cdot,\xi_{{\mathbf{g}}})$.
@@
The nonsmooth expectation constraints in Eq. P pose significant challenges for designing efficient and reliable stochastic methods, as the projection onto an expectation-constraint set is prohibitively expensive. To circumvent the costly projection step, several methods such as online convex method [63], augmented Lagrangian method (ALM) [38], inexact proximal point method (IPP) [10, 28, 46] and switching subgradient method (SSG) [27, 33] have been explored. Among these methods, the online convex method, ALM, and IPP all require at least an inner subsolver to address their respective subproblems. The online convex method [63] is developed specifically for problems with convex constraints. With nonconvex constraints and under suitable regularity conditions, Li et al. [38] and Xiao et al. [58] demonstrate the effectiveness of the ALM applied to problem Eq. P, yielding competitive numerical results. Xiao et al. [58] establishes convergence to a KKT point. To derive complexity results to an $(\epsilon,\epsilon)$-KKT point (see Definition 1.3), Li et al. [38] assume additional smoothness on $\mathbf{g}$} and exploit this structure to efficiently solve each nonconvex ALM subproblem by an accelerated method. Under a certain constraint qualification (CQ), Boob et al. [10] and Ma et al. [46] respectively provide convergence results and complexity analysis for stochastic IPP methods to find an $(\epsilon,\epsilon)$-KKT point. These methods iteratively solve strongly convex subproblems by incorporating a proximal term into the objective and constraint functions. Notably, the complexity results in [10] represent the state-of-the-art results for problem Eq. P under nonsmooth weakly convex settings.
@@
The aforementioned methods utilize optimal subroutines for solving inner subproblems to achieve low-order complexity results. Implementing them needs extra efforts to design appropriate checkable stopping criteria for inner subsolvers or requires careful tuning of the inner iteration counts to reach a satisfactory empirical performance. This disadvantage is especially pronounced in stochastic settings, where the randomness complicates the verification of a stopping condition and efficient tuning. Therefore, algorithms that avoid solving subproblems by an iterative subroutine and instead update the solution directly at each iteration are often more desirable for solving expectation-constrained problems in practice. However, such methods are still limited. Both [2, 54] design FOMs based on penalty or ALM for constrained stochastic programming but they require smoothness of the objective (possibly with a proximable nonsmooth convex term) and constraint functions. Huang and Lin [27] apply SSG [52] to problem Eq. P. The SSG is easier to implement while delivering competitive numerical performance. Under a (uniform) Slater-type CQ (see 2), the SSG adapts its stepsize depending on whether the current iterate is near feasible or not, switching between fixed and adaptive steps based on Polyak’s rules [52]. This adaptation is crucial in maintaining near feasibility to achieve a near KKT point. However, the SSG requires frequent checks for near-feasibility at generated iterates, resulting in a higher CFC and overall computational complexity compared to the IPP method in [10] when solving the stochastic nonconvex problem Eq. P. Furthermore, in the stochastic setting, Huang et al. [27] establish convergence guarantees only under the assumptions that $\mathbf{g}$} is convex and satisfies a light-tail condition.
@@
Without the need for feasibility checking, we propose a simple SPIDER-type Stochastic Subgradient FOM for solving the Expectation constrained problem Eq. P, named 3S-Econ. At the $k$-th iteration, 3S-Econ requires unbiased stochastic subgradient $\boldsymbol{\zeta}_{f}^{(k)}$ of $f$ and $\boldsymbol{\zeta}_{g_{i}}^{(k)}$ of $g_{i}$, evaluated at the iterate ${\mathbf{x}}^{(k)}$, along with an approximation $u_{i}^{(k)}$ of $g_{i}({\mathbf{x}}^{(k)})$, for all $i\in[m]$. Using this information, it performs the following update step:
@@
(2) ${\mathbf{x}}^{(k+1)}=\mathrm{Proj}_{\mathcal{X}}\left({\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{f}^{(k)}+\beta\sum_{i=1}^{m}\mathrm{Proj}_{[0,1]}\left(\textstyle\frac{u_{i}^{(k)}}{\nu}\right)\boldsymbol{\zeta}_{g_{i}}^{(k)}\right)\right),$
@@
where $\beta>0$ is a penalty parameter, $\nu>0$ is a smoothing parameter, and $\alpha_{k}>0$ denotes the step size at iteration $k$. We will show that 3S-Econ is not only easy to implement but also enjoys lower-order complexity than the best-known results to produce an $(\epsilon,\epsilon)$-KKT point of problem Eq. P; see Table 1.
@@ 1.1 Applications
Neyman-Pearson classification. Consider a classification problem involving training data partitioned into $M$ classes, denoted by $\mathcal{D}_{i}\subseteq\mathbb{R}^{d}$ for $i\in[M]$. Let $\phi$ be a loss function, $\{{\mathbf{x}}_{i}\}_{i\in[M]}$ denote model variables, $D>0$, and $\kappa$ be a prescribed threshold. The multi-class Neyman-Pearson classification aims to minimize the average pairwise loss for a specified class (e.g., class ${\mathcal{D}}_{1}$), while controlling the average pairwise losses for all other classes below a threshold. Formally, the optimization problem is given by
(3) $\min_{\left\|\mathbf{x}_{i}\right\|\leq D,i\in[M]}\frac{1}{\left|\mathcal{D}_{1}\right|}\sum_{l>1}\sum_{{\mathbf{a}}_{1}\in\mathcal{D}_{1}}\phi\left(\mathbf{x}_{1}^{\top}{\mathbf{a}}_{1}-\mathbf{x}_{l}^{\top}{\mathbf{a}}_{1}\right),$
$\text{ s.t. }\frac{1}{\left|\mathcal{D}_{i}\right|}\sum^{M}_{l=2,l\neq i}\sum_{{\mathbf{a}}_{i}\in\mathcal{D}_{i}}\phi\left(\mathbf{x}_{i}^{\top}{\mathbf{a}}_{i}-\mathbf{x}_{l}^{\top}{\mathbf{a}}_{i}\right)\leq\kappa,\quad i=2,3,\ldots,M.$
Clearly, Eq. 3 is a special instance of Eq. P with $f$ having a finite-sum structure, and $\mathbb{P}_{{\mathbf{g}}}$ corresponds to a uniform distribution over sample sets associated with $\{{\mathcal{D}}_{i}\}_{i=2,\ldots,M}$.
@@
Fairness-constrained classification. This application aims to ensure that classification models treat different groups equitably, avoiding biased outcomes. Traditional models focus on maximizing accuracy, which can lead to unfair performance across different groups of data, such as males and females. Fairness constraints are applied to balance outcomes for all groups, using principles like demographic parity [1, 35] or equalized odds [26]. Achieving fairness requires trade-offs, as improving fairness can reduce overall accuracy. This research area is particularly important in fields like healthcare and criminal justice [17].
Let us consider a classification problem with receiver operating characteristic (ROC)-based fairness [56]. Suppose the data are partitioned into three parts: a dataset $\mathcal{D}=\left\{(\mathbf{a}_{i},b_{i})\right\}_{i=1}^{n}$, and two labeled subgroups $\mathcal{D}_{\mathrm{p}}=\left\{\left(\mathbf{a}^{\mathrm{p}}_{i},b^{\mathrm{p}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{p}}}$ and $\mathcal{D}_{\mathrm{u}}=\left\{\left(\mathbf{a}^{\mathrm{u}}_{i},b^{\mathrm{u}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{u}}}$, defined according to a binary protected attribute (e.g., race or gender). Here, group $\mathcal{D}_{\mathrm{p}}$ corresponds to the minority subgroup with historically less favorable treatment, while group $\mathcal{D}_{\mathrm{u}}$ represents the majority subgroup. The fairness-constrained problem can be formulated as
(4) $\vskip-2.84526pt\min_{{\mathbf{x}}:\|{\mathbf{x}}\|\leq D}\Phi_{\Theta}({\mathbf{x}}),\,\,\text{{s.t.}}\,\,\Phi({\mathbf{x}})\leq\Phi^{*}+\kappa_{1},$